# Who is driving the conversation? Analysing the nodality of British MPs and journalists on Twitter

### v1

@@ Abstract
Who sets the policy agenda? Much political science research has been dedicated to investigating the influences at work in policy processes. This paper looks at the actors influencing the policy agenda through the lens of “nodality” —the capacity to share information and to be at the centre of social and information networks. Nodality has traditionally been seen as a capacity or policy tool relating to government as a set of institutions. But in a digital world, there is greater possibility for individual actors using social media to develop their own nodality and to shape the policy conversation. We analyse information networks on Twitter to quantify the nodality of two sets of actors in the UK political system - Members of Parliament (MPs) and accredited journalists – in conversations on four policy topics: The Russia-Ukraine War, the Cost-of-Living Crisis, Brexit and COVID-19. We perform a Transfer Entropy analysis and a Principle Component Analysis (PCA) using different network centralities while ensuring reproducibility and interpretability. We define and measure two dimensions of nodality; inherent nodality, associated with an actor’s institutional position, transferable between topics, and individual nodality, measuring active engagement in a policy conversation on a specific topic. These two dimensions are validated through a linear model showing how the influence on a particular topic exercised by an actor stems from the sum of nodality scores across these two dimensions. We thereby show the relative influence of different types of actor on policy conversations. More generally, we provide a methodology for measuring individual and institutional drivers of actor influence, and understanding where influence relates to individual or institutional characteristics.
@@ 1 Introduction
Framed within the context of policy discourse in the UK, we seek to analyse the nodality of different actors on online platforms to determine how they drive the competition to shape policy conversations. While our ultimate aim is to analyse the relative nodality of a more diverse set of actors, including governmental departments, NGOs, local councils and public figures, in this paper we first test our methodology on a smaller subset of elite actors in the UK: Members of Parliament (MPs) and accredited journalists. We propose a new method to dynamically measure nodality on social media, where information spreads widely and interactions are instantaneous. Although our methodology here is applied to Twitter (now X), it can be extended to other platforms and across a wide range of actors. Similar works include the study of exposure to misinformation from political elites in the USA [Mosleh and Rand, 2022], or the spread of news of COVID-19 from sources with different reputations [Gravino et al., 2022].
@@
Influence on social media has been researched extensively in the field of communications (see for example, [Dubois and Gaffney, 2014, Anger and Kittl, 2011, Bakshy et al., 2011, Mei et al., 2015, Riquelme and González-Cantergiani, 2016, Esteve Del Valle and Borge Bravo, 2018, Rezaie et al., 2020, Zhang et al., 2023]). On X (previously Twitter), as with other social media sites, we assume that an actor has an influence on a particular topic when their activity generates activity by others (tweeting, retweeting, liking, replying or quoting) on the same topic. Understanding actors’ influence on the policy and media agenda will in part relate to these dynamics in policy conversations, which can be studied using transfer entropy, thus allowing for temporal and predictive analysis [Schreiber, 2000, James et al., 2016, Shorten et al., 2021].
@@
In this paper, we investigate the nodality of two types of actors – MPs and journalists – that both play key roles in information exchange and influence in a policy system. Political representatives have an obligation to feed the concerns of their constituents into policy debates, while journalists disseminate information from policy discussions out to the public and wider information environment. We are able to obtain data for a substantial proportion of both sets of actors, which includes 581 out of 650 MPs and 8606 out of 9575 journalists [^1] The missing users do not have a Twitter account or the handles given could not be found. The dataset was also limited by the Twitter API policy which limits the number of tweets that can be retrieved per user to 3200.. This approach helps to address a methodological challenge faced by previous work on policy actors and social media, where the differential attention that actors receive relate to both institutional and individual-level characteristics. For institutional differences, MPs can hold ministerial or shadow ministerial positions, which means that they are more likely than backbench MPs to receive attention (through more followers, for example), while journalists working for mainstream news channels or newspapers are likely to receive more attention by virtue of their institutional homes, while independent journalists may struggle to receive attention. There are institutional constraints for both sets of actors on social media: ministers often need to follow a Government or party line, and journalists may need to follow guidelines of their publishers. For individual characteristics, some individual MPs may become active and influential on a certain topic, possibly by virtue of their constituency being differently affected by an issue. Individual journalists may acquire a social media presence over and above the reputation of the publications for which they work. Our approach allows us to develop measures of nodality across these ’institutional’ and ’individual’ dimensions, thereby understanding where influence relates to an actor’s institutional position, and where it relates to individual-level activity, reputation or characteristics. Developing this methodology may help in the analysis of social media activity more generally, such as understanding widely varying levels of hate speech and abuse directed at representatives [Vidgen et al., 2020].
@@
We test the hypothesis that that influence will vary across different types of actor and policy topics on social media platforms [Bright, 2016, Bright et al., 2020, Sprejer et al., 2022, Parmelee, 2013, Usherwood and Wright, 2017]. We do this by using data from social media conversations around four policy topics in the year 2022: (1) the Russia-Ukraine war (2) the challenging economic situation facing ordinary citizens (cost-of-living crisis) (3) Britain’s departure from the EU (Brexit) and (4) the aftermath of the COVID-19 pandemic. All four issues were contested in policy terms and related to an increasingly challenging economic situation, driven by Brexit [Diamond, 2023, Dhingra and Sampson, 2022], the pandemic [Blundell et al., 2022], the Russia-Ukraine war [Mbah and Wasum, 2022, Liadze et al., 2023] and the resulting energy crisis [Harari et al., 2022, Francis-Devine et al., 2022], as well as a “mini-budget” in the autumn which led to sharp and sustained rises in interest rates [Chadha, 2022, Tosun and Lucey, 2023, Li, 2023]. We analyse Twitter data obtained for 581 MPs and 8,606 journalists registered in the UK between January 14, 2022 and January 13, 2023. We classify tweets using a weak-supervision classifier [Ratner et al., 2019, Castro-Gonzalez et al., 2024] and construct a weighted directed graph for each topic, where actors are nodes, and any interaction is captured as an information edge. To obtain group-level influence measures, we employ Transfer Entropy analysis on their activity time series, while inherent and active nodalities as defined above are derived through a Principal Component Analysis (PCA). We hypothesise that we will be able to distinguish institutional nodality as transferable across various topics, while active nodality is specific to a particular topic.
@@ 4.5 Relation between influence score and active and inherent nodality
Eigenvalue centrality
A high eigenvalue centrality is related to the probability that information will pass through a particular node. In this sense, a node with a respective high eigenvalue centrality is interpreted as a node whose activity will most probably be seen in the conversation network by the other actors.
Visibility
Betweenness centrality
Betweenness centrality is related to the likelihood that information will pass through a particular node. Thus, an actor who is important to deliver a message in the conversation network is interpreted to be one with a high betweenness centrality.
Within-nodes importance
Strength
The strength of a node takes into account the number of links with other nodes and the weights associated with them. Thus, a node with a small number of ties but with high weights can be as “strong” as one with a high number of ties with low weights.
Engagement
Authority (HITS algorithm)
A high authority score reflects, in our case, a node that has a high number of out-degree links, thus having a high number of other nodes interacting with it. The algorithm reflects in this sense the authority of the node in the network [Kleinberg, 1999].
Authority
Hubs (HITS algorithm)
A high hubs score reflects, in contrast with the authority score, a node that has a high number of in-degree links, thus reflecting the node interacting with a high number of other nodes. The algorithm reflects in this sense how much the node acts as a hub of information within the network [Kleinberg, 1999].
Information transmission
Funnel bandwidth $\nu$
This term reflects how much amplification can a node give to other’s messages when interacting with them. By having an important number of followers, a node can give others visibility by interacting with them. The more interactions, the more visibility a node gives to others. The former is then acting as a funnel of information
Information transmission
Amplification bandwidth $\mu$
The amplification bandwidth relates to who interacts with a given node. The more nodes interacting with you, the more frequent those interactions are, and the more followers your interlocutors have, the more amplified your activity will be.
Amplification
@@ Acknowledgments.
Our code and experimental results are available in our GitHub repository: https://github.com/LeonardoCastro/Nodality.

### v2

@@ Abstract
Who sets the policy agenda? In this paper, we explore the roles of policy actors in agenda setting by studying their relative influence in policy-related discussions. Our approach builds on “nodality” —a concept in political science that determines the capacity of an actor to share information and to be at the centre of information networks. We propose a novel methodology that quantifies the nodality of all individual actors in any conversation by analysing a comprehensive set of their centrality measures in the related information network. We combine this with the analysis of the activity time-series, of the related conversation (or topic), to demonstrate how nodality scores relate to the capacity to drive topic-related activity. Here we analyse policy-related discussions on X (previously Twitter) and quantify the nodality of two sets of actors in the UK political system —Members of Parliament (MPs) and accredited journalists - on four policy topics: The Russia-Ukraine War, the Cost-of-Living Crisis, Brexit and COVID-19. Our results show that the capacity to influence the activity related to a topic is significantly and positively associated with nodality. In particular, we identify two dimensions of nodality that drive the capacity to influence topic-related activity. The first is “active nodality”, which reflects the level of topic-related engagement an individual actor has on the platform. The second dimension is “inherent nodality” which is entirely independent of the platform and reflects the actor’s institutional position (such as an MP in a front-bench role, or a journalist’s position at a prominent media outlet).
@@ 1 Introduction
Framed within the context of policy discourse in the UK, we seek to analyse the nodality of different actors on online platforms to determine how they drive the competition to shape policy conversations. While our ultimate aim is to analyse the relative nodality of a more diverse set of actors, including governmental departments, NGOs, local councils and public figures, in this paper we first test our methodology on a subset of elite actors in the UK: Members of Parliament (MPs) and accredited journalists. We propose a new method to dynamically measure nodality on social media, where information spreads widely and interactions are instantaneous. Although our methodology here is applied to Twitter (now X), it can be extended to other platforms and across a wide range of actors. Similar works include the study of exposure to misinformation from political elites in the USA [Mosleh and Rand, 2022], or the spread of news of COVID-19 from sources with different reputations [Gravino et al., 2022].
@@
Influence on social media has been researched extensively in the field of communications (see for example, [Dubois and Gaffney, 2014, Anger and Kittl, 2011, Bakshy et al., 2011, Mei et al., 2015, Riquelme and González-Cantergiani, 2016, Esteve Del Valle and Borge Bravo, 2018, Rezaie et al., 2020, Kolic et al., 2022, Flamino et al., 2023, Zhang et al., 2023]). On X (previously Twitter), as with other social media sites, we assume that an actor has an influence on a particular topic when their activity generates activity by others (tweeting, retweeting, liking, replying or quoting) on the same topic. Understanding actors’ influence on the policy and media agenda will in part relate to these dynamics in policy conversations, which can be studied using transfer entropy, thus allowing for temporal and predictive analysis [Schreiber, 2000, James et al., 2016, Shorten et al., 2021].
@@
In this paper, we investigate the nodality of two types of actors – MPs and journalists – that both play key roles in information exchange and influence in a policy system. Political representatives have an obligation to feed the concerns of their constituents into policy debates, while journalists disseminate information from policy discussions out to the public and wider information environment. We are able to obtain data for a substantial proportion of both sets of actors, which includes 581 out of 650 MPs and 8606 out of 9575 journalists [^1] The missing users do not have a Twitter account or the handles given could not be found. The dataset was also limited by the Twitter API policy which limits the number of tweets that can be retrieved per user to around 3200.. This approach helps to address a methodological challenge faced by previous work on policy actors and social media, where the differential attention that actors receive relate to both institutional and individual-level characteristics. For institutional differences, MPs can hold ministerial or shadow ministerial positions, which means that they are more likely than backbench MPs to receive attention (through more followers, for example), while journalists working for mainstream news channels or newspapers are likely to receive more attention by virtue of their institutional homes, while independent journalists may struggle to receive attention. There are institutional constraints for both sets of actors on social media: ministers often need to follow a Government or party line, and journalists may need to follow guidelines of their publishers. For individual characteristics, some individual MPs may become active and influential on a certain topic, possibly by virtue of their constituency being differently affected by an issue. Individual journalists may acquire a social media presence over and above the reputation of the publications for which they work. Our approach allows us to develop measures of nodality across these ’institutional’ and ’individual’ dimensions, thereby understanding where influence relates to an actor’s institutional position, and where it relates to individual-level activity, reputation or characteristics. Developing this methodology may help in the analysis of social media activity more generally, such as understanding widely varying levels of hate speech and abuse directed at representatives [Vidgen et al., 2020].
@@
We test the hypothesis that that influence will vary across different types of actor and policy topics on social media platforms [Bright, 2016, Bright et al., 2020, Sprejer et al., 2022, Parmelee, 2013, Usherwood and Wright, 2017]. We do this by using data from social media conversations around four policy topics in the year 2022: (1) the Russia-Ukraine war (2) the challenging economic situation facing ordinary citizens in the UK (Cost-of-Living crisis) (3) Britain’s departure from the EU (Brexit) and (4) the aftermath of the COVID-19 pandemic. All four issues were contested in policy terms and related to an increasingly challenging economic situation, driven by Brexit [Diamond, 2023, Dhingra and Sampson, 2022], the pandemic [Blundell et al., 2022], the Russia-Ukraine war [Mbah and Wasum, 2022, Liadze et al., 2023] and the resulting energy crisis [Harari et al., 2022, Francis-Devine et al., 2022], as well as a “mini-budget” in the autumn which led to sharp and sustained rises in interest rates [Chadha, 2022, Tosun and Lucey, 2023, Li, 2023]. We analyse Twitter data obtained for 581 MPs and 8,606 journalists registered in the UK between January 14, 2022 and January 13, 2023. We classify tweets using a weak-supervision classifier [Ratner et al., 2019, Castro-Gonzalez et al., 2024] and construct a weighted directed graph for each topic, where actors are nodes, and any interaction is captured as an information edge. To obtain group-level influence measures, we employ Transfer Entropy analysis on their activity time series, while inherent and active nodalities as defined above are derived through a Principal Component Analysis (PCA). We hypothesise that we will be able to distinguish institutional nodality as transferable across various topics, while active nodality is specific to a particular topic.
@@ 4.5 Relation between influence score and active and inherent nodality
Eigenvalue centrality A high eigenvalue centrality is related to the probability that information will pass through a particular node. In this sense, a node with a respective high eigenvalue centrality is interpreted as a node whose activity will most probably be seen in the conversation network by the other actors. Visibility
Betweenness centrality Betweenness centrality is related to the likelihood that information will pass through a particular node. Thus, an actor who is important to deliver a message in the conversation network is interpreted to be one with a high betweenness centrality. Within-nodes importance
Strength The strength of a node takes into account the number of links with other nodes and the weights associated with them. Thus, a node with a small number of ties but with high weights can be as “strong” as one with a high number of ties with low weights. Engagement
Authority (HITS algorithm) A high authority score reflects, in our case, a node that has a high number of out-degree links, thus having a high number of other nodes interacting with it. The algorithm reflects in this sense the authority of the node in the network [Kleinberg, 1999]. Authority
Hubs (HITS algorithm) A high hubs score reflects, in contrast with the authority score, a node that has a high number of in-degree links, thus reflecting the node interacting with a high number of other nodes. The algorithm reflects in this sense how much the node acts as a hub of information within the network [Kleinberg, 1999]. Information transmission
Funnel bandwidth $\nu$ This term reflects how much amplification can a node give to other’s messages when interacting with them. By having an important number of followers, a node can give others visibility by interacting with them. The more interactions, the more visibility a node gives to others. The former is then acting as a funnel of information Information transmission
Amplification bandwidth $\mu$ The amplification bandwidth relates to who interacts with a given node. The more nodes interacting with you, the more frequent those interactions are, and the more followers your interlocutors have, the more amplified your activity will be. Amplification
@@ Acknowledgments.
Twitter data is made available in accordance with Twitter’s terms of service. Tweet IDs and analysis code are available at https://github.com/Turing-Online-Safety-Codebase/nodality. The corresponding tweets can be downloaded using the official Twitter API (https://developer.twitter.com/en/docs/twitter-api)